OpenStudy (anonymous):

The half-life of strontium 90 is 28 years. (a) Obtain an exponential decay model for strontium 90 in the form Q(t) = Q0e−kt.

OpenStudy (anonymous):

Use your model to predict, to the nearest year, the time it takes two-thirds of a sample of strontium 90 to decay.

OpenStudy (campbell_st):

well you need to find the value of k so let the initial quantity be 1 so the quantity after 28 years is 1/2 so substituting into the formula \[0.5 = 1 \times e^{-k \times 28}\] take the log of both sides \[\log_{e}(0.5) = -28k\] you should be able to evalaute k...from here..

OpenStudy (anonymous):

k=.01075

OpenStudy (campbell_st):

ummm not quite \[k = \frac{\log_{e}(0.5)}{-28} = \] i got 0.024755...